# LazyEviction: Lagged KV Eviction with Attention Pattern Observation for Efficient Long Reasoning

### v1

@@ Abstract
Large Language Models (LLMs) exhibit enhanced reasoning capabilities by employing Chain-of-Thought (CoT). However, the extended reasoning sequences introduce significant GPU memory overhead due to increased key-value (KV) cache size, particularly in tasks requiring long reasoning sequences, such as mathematics and programming. Existing KV cache compression methods mitigate memory bottlenecks but struggle in long reasoning tasks. In this paper, we analyze attention patterns in reasoning tasks and reveal a Token Importance Recurrence phenomenon: a large proportion of tokens receive renewed attention after multiple decoding steps, which is failed to capture by existing works and may lead to unpredictable eviction on such periodically critical tokens. To address this, we propose LazyEviction, a lagged KV eviction framework designed to maintain reasoning performance while reducing KV memory. LazyEviction is an Observation Window-based Lagged Eviction Mechanism retaining latent recurring tokens by performing lagged evictions across decoding steps, which contains two key components: (1) Recurrence Interval Tracking for capturing temporal variations in token importance, and (2) an Maximum Recurrence Interval-Centric Eviction Policy that prioritizes eviction based on tokens’ recurrence patterns. Extensive experiments demonstrate that LazyEviction reduces KV cache size by 50% while maintaining comparable accuracy on mathematics reasoning datasets, outperforming state-of-the-art methods. Our findings highlight the importance of preserving recurring tokens, which are critical for maintaining knowledge continuity in multi-step reasoning tasks.
@@ 1 Introduction
Large Language Models (LLMs) have emerged with advanced capabilities in complex reasoning tasks by enabling Chain-of-Thought (CoT) to elicit step-by-step inference wei2022chain ; kojima2022large . Recent advancements, such as OpenAI’s o1 jaech2024openai and DeepSeekR1 guo2025deepseek , further demonstrate that scaling up CoT lengths from hundreds to thousands of reasoning steps could continuously improve LLM reasoning. However, the increased reasoning sequences introduce substantial memory and computational overhead. Due to the autoregressive nature of LLM decoding, longer CoT outputs lead to proportional increases in GPU memory for caching key-value(KV) states. These issues become particularly pronounced when reasoning sequences extend into thousands of reasoning tokens, resulting in significant GPU memory costs. For example, in mathematics cobbe2021training ; hendrycks2021measuring and programming tasks chen2021evaluating , the reasoning sequences can grow up to 16k tokens, resulting in more than 100GB KV cache with a batch size of 32, which exceeds the memory capacity of even high-end GPUs.
KV cache compression has emerged as a promising approach to alleviate the memory bottleneck caused by increasing KV cache sizes. Unlike traditional long-context input tasks that focus on the prefilling phase li2024snapkv ; cai2024pyramidkv ; feng2024ada ; fu2024not , long-reasoning tasks perform long-context generation that requires dynamic KV cache compression in multiple decoding steps. The critical challenge lies in compressing generated KV caches while minimizing performance degradation. Existing approaches mitigate this through selective eviction strategies: early works liu2023scissorhands ; xiao2023efficient preserve subsets of recent and initial KV pairs, while others oren2024transformers ; chen2024nacl ; he2025treekv ; zhang2023h2o ; adnan2024keyformer leverage attention scores to prioritize critical tokens and evict others, which can be further divided into two types: (a) Current Attention-based Eviction (TOVA oren2024transformers , NACL chen2024nacl , TreeKV he2025treekv ) and (b) Cumulative Attention based Eviction (H2O zhang2023h2o , Keyformer adnan2024keyformer , MorphKV ghadia2025dialogue , RaaS hu2025efficient ). However, as shown in Fig. 1, these works are greedy eviction strategies at each decoding step, which discards tokens deemed temporarily unimportant while overlooking their potential future importance. We empirically observe that recurrence of token importance is a frequent phenomenon in multi-step reasoning processes, which typically involve conditional or summary information within reasoning chains, aligning with the capabilities in verification, backtracking, and summarization in reasoning models gandhi2025cognitive . The previous works approaches risk prematurely discarding crucial tokens during reasoning steps, ultimately sacrificing task accuracy.
To achieve efficient KV cache compression while preserving reasoning performance, we analyze attention patterns during decoding phases. Our investigation reveals a Token Importance Recurrence phenomenon: a large proportion of tokens in reasoning tasks receive renewed attention after multiple decoding steps. In this paper, we refer to such tokens as recurring tokens, where the attention score of recurring tokens may be low within an interval while suddenly increasing in the future. Empirical analysis demonstrates that these tokens often contain information vital for later reasoning stages, and premature eviction of recurring tokens leads to catastrophic performance degradation. Crucially, existing KV compression methods usually fail to identify such tokens due to inherent limitations.
@@
Figure 1: Comparison of different KV Eviction methods. The dark squares represent that the token has a relatively higher attention score. (a) Current Attention-based Eviction execute stepwise evictions using immediate attention scores. (b) Cumulative Attention-based Eviction integrate historical attention for eviction decisions. Both (a) and (b) fails to preserve recurring tokens during their low-attention intervals. (c) LazyEviction performs lagged KV evictions based on the observation window to detect latent recurring tokens and prevent prematurely discarding them.
@@
To address this challenge, we propose LazyEviction, a novel Observation Window-based Lagged Eviction Mechanism. Instead of per-step greedy decisions, we perform lagged KV evictions at every $W$ decoding steps to detect latent recurring tokens and prevent prematurely discarding them. There are two key componebts: (1) Recurrence Interval Tracking: For each token, we record its Maximum Recurrence Interval (MRI)—the longest observed period between its attention spikes during decoding to capture the temporal variation of token importance. (2) MRI-Centric Eviction Policy: When reaching memory budgets, we prioritize evicting tokens whose $\Delta T$ exceeds their MRI, where $\Delta T$ is the time elapsed since the token’s last important recurrence. Tokens with $\Delta T<MRI$ are retained as they still have the potential to become important in the future. LazyEviction shifts from step-wised greedy eviction to window-wised predictive retention: by continuously tracking MRI and executing lagged eviction during observation windows, we can effectively detect latent recurring tokens and prevent prematurely discarding them. To sum up, our key contributions are:
  1. 1.
@@
    We observe Token Importance Recurrence in the reasoning task, where specific tokens exhibit recurrent attention patterns (i.e., being re-activated after multiple decoding steps). These recurring tokens often represent foundational formulations or intermediate conclusions which are crucial for maintaining knowledge continuity in multi-step reasoning.
  2. 2.
@@
    We propose the LazyEviction, a Lagged KV Eviction scheme integrating temporal observation window and recurrence interval tracking. Our framework enables prediction of future token importance, thereby preventing premature eviction of recurring tokens.
  3. 3.
@@
    We have implemented LazyEviction on the latest large reasoning model (e.g. DeepSeek-R1-Distill-Qwen and DeepSeek-R1-Distill-Llama). It has been verified on mathematics reasoning datasets that our method can maintain comparable performance while reducing 50% KV budget, surpassing the SOTA KV cache compression methods.
@@ 2 Related Works
Various approaches have been developed to improve LLM’s efficiency in handling long reasoning tasks. These methods can be classified into three main categories: KV cache compression, long reasoning compression, and system-level optimizations.
KV cache compression: Considering the inherent sparsity of attention mechanisms, these works focus on retaining critical tokens’ KV pairs while evicting non-essential ones within constrained memory budgets. Recent works focus on compressing KV cache during the model decoding phase. We summarize the existing works and analyze their limitations as follows:
  1. 1.
    Static Positional Retention: Methods like Scissorhands liu2023scissorhands and StreamingLLM xiao2023efficient employ static retention of initial and recent tokens under fixed budgets. While sustaining text coherence, their rigid retention policies cannot detect latent recurring tokens.
  2. 2.
    Current Attention-based Eviction: TOVA oren2024transformers , NACL chen2024nacl , and TreeKV he2025treekv execute stepwise evictions using immediate attention scores. This myopic strategy fails to preserve recurring tokens during their low-attention intervals.
  3. 3.
    Cumulative Attention-based Eviction: H2O zhang2023h2o , Keyformer adnan2024keyformer , and MorphKV ghadia2025dialogue integrate historical attention for eviction decisions. RaaS hu2025efficient extends this through dynamic updated timestamp. These approaches still mistakenly regard latent recurring tokens as unimportant due to continuous low historical scores.
Long Reasoning Compression: Recent works have focused on directly compressing the output of reasoning models. For instance, TALE han2024token and SoT aytes2025sketch adopts prompts engineering to control the length of CoT outputs. TokenSkip xia2025tokenskip achieves output compression through fine-tuning on condensed CoT datasets. InftyThink yan2025inftythink and LightThinker zhang2025lightthinker equip models with summarization capabilities to reduce token consumption per reasoning step. Approaches like COCONUT hao2022training , CCoT cheng2024compressed , and SoftCoT xu2025softcot enable reasoning directly in compressed latent spaces. Distinctively, LazyEviction maintains the original output length of LLMs while enhancing memory and computational efficiency through optimized KV cache usage, demonstrating compatibility with these compression approaches.
System-level Optimizations: By optimizing GPU memory usage and batch processing, vLLM kwon2023efficient , FlashAttention dao2022flashattention , and FlashDecoding hong2023flashdecoding imporve both time and memory efficiency in decoding. Additionally, approaches like QUEST tang2024quest and OmniKV hao2025omnikv store all KV caches in CPU and dynamically schedule them during decoding phase. Since these methods retain all KV caches, they can handle recurring tokens proposed in this paper. However, they incur non-negligible I/O latency from CPU-GPU transfers and rely on precise critical cache selection. By offloading non-critical tokens to CPU, LazyEviction can be compatible with such techniques.
@@ 3 Observations and Motivations
Figure 2: (a) shows the performance degradation for SOTA Methods on reasoning tasks. With the same KV cache compression ratio (e.g., 50%), compared with traditional language modeling tasks dataset PG-19 rae2019compressive , the performance of both H2O and TOVA has decreased by more than 20% on GSM8K dataset. (b) is the visualization of the importance variation by selecting Top-50% important tokens. Tokens at the same position show different importance at different decoding steps.
Figure 3: Visualization of Token Importance Recurrence. We observe attention maps across different heads of DeepSeek-R1-Distill-Llama-8B. We find most attention maps (>80%) show Token Importance Recurrence. (a) and (b) show an attention map with recurring tokens and their corresponding positions. (c) statistically analyzed the MRI distribution in different models among different tasks.
The goal of KV cache eviction is to identify and evict KV pairs that are non-critical for future decoding steps. To preserve the correctness of subsequent decoding results, it is critical to retain tokens with potential future importance. In this section, we first analyze the previous works and observe attention patterns in reasoning tasks. We summarize our findings as follows:
Finding 1: Previous KV cache compression algorithm, while demonstrating high performance in general long-context generation scenarios, exhibit degraded performance when applied to long reasoning tasks.
We compare the performance of H2O zhang2023h2o and TOVA oren2024transformers on reasoning tasks versus general long-context generation tasks in Fig.2(a), H2O and TOVA report competitive performance on standard language modeling dataset, but exhibit significant performance degradation on GSM8K under the same KV cache compression ratio. To diagnose this failure, we track the positions of top-50% important tokens across decoding steps in Fig.2(b) and observe that tokens critical in later steps are often missing in earlier steps. Due to their low importance scores in intermediate steps, these tokens are prone to premature eviction in prior works.
Finding 2: Token Importance Recurrence A prevalent phenomenon in reasoning tasks, where tokens initially exhibit strong attention weights, receive low attention weights during later decoding steps, and subsequently regain attention after a certain decoding steps interval. Recurring tokens contain critical information for the reasoning and prematurely evicting them will cause knowledge in-continuity in multi-step reasoning tasks.
@@
Finding 3 Most recurring tokens exhibit an MRI far smaller than the generated length and can be detected via an observation window, which motivates our design of the Observation Window-based Lagged Eviction Mechanism.
As shown in in Fig. 3(c), We statistically analyzed the MRI distribution in different models among different tasks. As the output length of the model increases, the MRI gradually becomes larger, but most tokens’ MRI remian relatively small. For example, the output length of the Qwen model on the MATH500 dataset can reach 8k, but we have statistically found that the MRI of 80% of the tokens is less than 175.Therefore, as long as the window size exceeds the MRI of most tokens, the majority of recurring tokens will be identified.
@@ 4 LazyEviction
In this section, we introduce our LazyEviction, which is a novel framework of Observation Window-based Lagged Eviction. We first model LazyEviction as a multi-step dynamic optimization problem in Appendix A.1. Unlike existing works that perform KV eviction decisions at every decoding step, our LazyEviction adopts a window size $W$ as the interval for KV eviction decisions. As illustrated in Fig. 4(a), during decision-making, LazyEviction always retains the most recent $W$ KVs to preserve local coherence. When the number of KVs exceeds the predefined budget $B(B\gg W)$, LazyEviction executes KV eviction. The overall procedure of our LazyEviction is summarized in Appendix. A.3.
To select $B-W$ KVs from past caches to retain, LazyEviction tracks importance variations (i.e., MRI) of each token at every decoding step (Recurrence Interval Tracking). For eviction decisions, importance scores, which are defined to indicate tokens’ potential future importance, are calculated based on MRI. Consequently, LazyEviction selectively retains only the tokens in the KV cache most likely to be critical in the future (MRI-Centric Eviction).
Figure 4: Overview of our proposed LazyEviction Framework, where the dark squares represent that the token has a relatively higher attention score. (a) LazyEviction performs eviction decision at intervals of $W$ steps. The workflow contains two key operations: (b) Dynamic MRI Tracking according to updated important timestamp, and (c) MRI-Centric Scoring during decision phases, where tokens predicted to be critical for future steps are retained. (d) The MRI-Centric Score fundamentally predicts future token importance by analyzing historical patterns of importance variation (i.e., MRI and the time elapsed since their latest timestamp).
Recurrence Interval Tracking
To identify importance variations for recurring tokens, we employ timestamps to track the time of being important. Consistent with the definition in RaaS hu2025efficient , whenever a token receives an attention score exceeding a threshold $\alpha$ (e.g., $\alpha=0.0001$), its latest timestamp is updated to the value of current decoding step $t$. Thus, during the whole decoding phase, LazyEviction dynamically maintains a timestamp vector $\mathbf{TS}_{t}$ for all retained tokens, which records only the most recent activation time of tokens (illustrated in the leftmost figure in Fig. 4(b)).
However, only relying on $\mathbf{TS}_{t}$ fails to capture the temporal patterns of Importance Recurrence. To address this, we introduce the MRI to record the longest interval between two consecutive activations in history. For newly generated tokens, MRI is initialized to 0. The MRI value is updated as follows:
$\mathbf{MRI}_{t}=\max\{\mathbf{MRI}_{t-1},\mathbf{TS}_{t}-\mathbf{TS}_{t-1}\}$ (1)
MRI-Centric Eviction
When $\left|{\mathcal{S}_{t}}\right|$ exceeds $B$, KV eviction are triggered at periodic intervals $t=kW$ ($k\in\mathbf{N}^{+}$). Leveraging the tracked historical importance dynamics via $\mathbf{MRI}_{t-1}$ and $\mathbf{TS}_{t}$. We aim to predict each token’s future importance using an MRI-Centric Importance Score. As illustrated in Fig 4(d) the score should be designed under the following two intuitive heuristics:
@@
    H1: Given a token $i$, if time elapsed since its last activation ($t-T{S_{t}}[i]$) is greater than its $MR{I_{t}}[i]$, the possibility to be important in the future will decrease. The larger the ratio between $t-T{S_{t}}[i]$ and $MR{I_{t}}[i]$, the less likely token $i$ is important.
@@
    H2: Tokens with smaller MRI values have higher possibility to be important in the future, as frequently recurring tokens are deemed more critical. Notably, if $MR{I_{t}}[i]=0$, the token has never been activated since its generation (MRI is initialized as 0), resulting in an importance score of zero.
Based on the above intuitive heuristics, the MRI-Centric Importance Score can be formalized as:
@@
We define ${}\exp\{\cdot\}$ term as H1-Score, which reflects the likelihood of the token regaining importance within the next observation window. Define $1/{MR{I_{t}}[i]}$ as H2-Score to prioritizes tokens with smaller MRI values. For each eviction decision, we first compute $\mathbf{I}_{t}$ with Eq. 2. Then, we retain the most recent $W$ KV pairs and the $\rm{Top}(B-W)$ KV pairs with the highest scores.
Disscussion about the choice of $W$ and $\alpha$
In LazyEviction, the observation window with size $W$ serves to detect recurring tokens. However, since the most recent $W$-th tokens must be preserved, the setting of $W$ inherently introduces a trade-off between local and global information retention. An excessively large $W$ risks losing critical past tokens, while an overly small $W$ may fail to detect numerous recurring tokens. To resolve this dilemma, we employ the observation from Fig. 3(c), setting $W$ as the MRI threshold corresponding to the 80% tokens. This threshold can be statistically determined through offline analysis by randomly selecting 50 samples. Our experiments in Sec. 5.3 analyze how $W$ affects reasoning performance.
Since the selection of $\alpha$ influences the temporal distribution of token timestamps, both extremely large and small $\alpha$ hinder effective identification of tokens’ importance variation. As the methodology for determining $\alpha$ has been thoroughly discussed in RaaS hu2025efficient , we directly adopt their approach in LazyEviction. We analyze the effect of setting different $\alpha$ in Sec. 5.3.

### v3

@@ Abstract
Large Language Models (LLMs) exhibit enhanced capabilities by Chain-of-Thought reasoning. However, the extended reasoning sequences introduce significant GPU memory overhead due to increased key-value (KV) cache. Existing KV cache compression methods mitigate memory bottlenecks but struggle in long reasoning tasks. In this paper, we analyze attention patterns in reasoning tasks and reveal a Token Importance Recurrence phenomenon: a large proportion of tokens regain high attention after multiple decoding steps, which is failed to capture by existing works and may lead to unpredictable eviction on such periodically critical tokens. To address this, we propose LazyEviction, an observation window-based lagged eviction framework retaining latent recurring tokens by prioritized eviction based on tokens’ recurrence patterns. Extensive experiments demonstrate that LazyEviction reduces KV cache by 50%~70% while maintaining comparable accuracy, outperforming existing KV cache compression baselines. Our implementation code can be found at https://github.com/Halo-949/LazyEviction.
LazyEviction: Lagged KV Eviction with Attention Pattern Observation for Efficient Long Reasoning
Haoyue Zhang1, Hualei Zhang1, Xiaosong Ma2, Jie Zhang1, Song Guo1, 1HKUST, 2HK PolyU Correspondence:Jie Zhang, csejzhang@ust.hk
@@ 1 Introduction
Large Language Models (LLMs) have emerged advanced capabilities in complex reasoning tasks by enabling Chain-of-Thought (CoT) to elicit step-by-step inference [1; 2]. Recent advancements, such as OpenAI’s o1 [3] and DeepSeekR1 [4], further demonstrate that scaling up CoT lengths from hundreds to thousands of reasoning steps could continuously improve LLM reasoning. However, the increased reasoning sequences introduce substantial memory and computational overhead. Due to the autoregressive nature of LLM decoding, longer CoT outputs lead to proportional increases in GPU memory for caching key-value (KV) states. These issues become particularly pronounced when reasoning sequences extend into thousands of reasoning tokens, resulting in significant GPU memory costs. For example, in mathematics [5; 6] and programming tasks [7], the reasoning sequences can grow up to 16k tokens, resulting in more than 100GB KV cache with a batch size of 32, which exceeds the memory capacity of even high-end GPUs.
@@
Figure 1: Comparison of different KV Eviction methods. The dark squares represent that the token has a higher attention score. (a) Current Attention-based Eviction executes stepwise evictions using immediate attention scores. (b) Cumulative Attention-based Eviction integrates historical attention for eviction decisions. Both (a) and (b) fail to preserve recurring tokens during their low-attention intervals. (c) LazyEviction performs lagged KV evictions based on the observation window to detect latent recurring tokens and prevent prematurely discarding them.
KV cache compression has emerged as a promising approach to alleviate the memory bottleneck caused by increasing KV cache sizes. Unlike traditional long-context input tasks that focus on the prefilling phase [8; 9; 10; 11], long-reasoning tasks perform long-context generation that requires compression in multiple decoding steps. The critical challenge lies in compressing generated KV caches while minimizing performance degradation. Existing approaches mitigate this through selective eviction strategies: early work [12] preserves subsets of recent and initial KV pairs, while others [13; 14; 15; 16; 18] leverage attention scores to prioritize critical tokens and evict others, which can be further divided into two types: (a) Current Attention-based Eviction [13; 14; 15] and (b) Cumulative Attention based Eviction [16; 18; 20; 19]. However, as shown in Fig. 1, these works are greedy eviction strategies at each decoding step, which discard tokens deemed temporarily unimportant while overlooking their potential future importance.
We empirically analyze attention patterns during multi-step reasoning processes. Our investigation reveals a frequent Token Importance Recurrence (TIR) phenomenon: a large proportion of tokens in reasoning tasks receive renewed attention after multiple decoding steps. In this paper, we refer to such tokens as recurring tokens, where the attention score of recurring tokens may be low within an interval while suddenly increasing in the future. Empirical analysis demonstrates that these tokens typically involve conditional or summary information within reasoning chains, aligning with the capabilities in verification, backtracking, and summarization in reasoning models [36]. The recurring tokens often contain information vital for later reasoning stages, and premature eviction of recurring tokens leads to catastrophic performance degradation. Crucially, existing KV compression methods usually fail to identify such tokens due to inherent limitations, thus risking prematurely discarding crucial tokens during reasoning steps, ultimately sacrificing task accuracy.
@@
To address this challenge, we propose LazyEviction, a novel Observation Window-based Lagged Eviction Mechanism. Instead of per-step greedy decisions, we perform lagged KV evictions at every $W$ decoding steps to detect latent recurring tokens and prevent prematurely discarding them. There are two key components: (1) Recurrence Interval Tracking: we record each token’s Maximum Recurrence Interval (MRI) (i.e., the longest observed period between its attention spikes) to capture the temporal variation of token importance. (2) MRI-Centric Eviction Policy: we prioritize evicting tokens whose $\Delta T$ exceeds their MRI, where $\Delta T$ is the time elapsed since the token’s last important recurrence. Tokens with $\Delta T<MRI$ are retained as they still have the potential to become important in the future. LazyEviction shifts from step-wised greedy eviction to window-wised predictive retention: by continuously tracking MRI and executing lagged eviction during observation windows, we can effectively detect latent recurring tokens and prevent prematurely discarding them. To sum up, our key contributions are:
@@
  • We observe token importance recurrence in the reasoning task, where specific tokens exhibit recurrent attention patterns. These recurring tokens often represent foundational formulations or intermediate conclusions that are crucial for maintaining knowledge continuity in multi-step reasoning.
@@
  • We propose LazyEviction, an observation window-based lagged KV eviction scheme integrating recurrence interval tracking. Our framework enables prediction of future token importance, thereby preventing premature eviction of recurring tokens.
@@
  • We have implemented LazyEviction on the latest large reasoning models ranging from 4B to 32B. It has been verified on multiple reasoning datasets that our method can maintain comparable performance while reducing 50%~70% KV budget, surpassing the SOTA KV cache compression methods.
@@ 2 Related Works
Various approaches have been developed to improve LLM’s efficiency in handling long reasoning tasks, which can be classified into three main categories: KV cache compression, reasoning path compression, and system-level optimizations.
KV cache compression. Considering the inherent sparsity of attention mechanisms, these works focus on retaining critical tokens’ KV pairs while evicting non-essential ones within constrained memory budgets. We summarize the existing works as follows:
(1) Static-Position retention: StreamingLLM  [12] employs static retention of initial and recent tokens under fixed budgets. While sustaining text coherence, their rigid retention policies cannot detect latent recurring tokens.
(2) Current Attention-based Eviction: TOVA  [13], NACL [14], and TreeKV [15] execute stepwise evictions using immediate attention scores. This myopic strategy fails to preserve recurring tokens during low-attention intervals.
(3) Cumulative Attention-based Eviction: Scissorhands  [17], H2O [16], Keyformer [18], and MorphKV [20] integrate tokens’ historical attention values for eviction decisions. RaaS [19] extends this through dynamic updated timestamp. These approaches still mistakenly regard latent recurring tokens as unimportant due to continuous low historical scores.
Recent works R-KV [37] and RPC [38] utilize token similarity to identify and evict redundant tokens in long reasoning tasks. Instead, LazyEviction directly focuses on the attention patterns in reasoning path. Our goal is to use the compressed KV cache to closely approximate the full KV attention map, ensuring the theoretical interpretability of performance.
Long Reasoning Compression. Recent works have focused on compressing the output of reasoning models. For instance, TALE [24] and SoT [26] adopt prompt engineering to control the length of CoT outputs. TokenSkip [25] achieves output compression through fine-tuning on condensed CoT data. InftyThink [27] and LightThinker [28] equip models with summarization capabilities to reduce tokens during the reasoning step. Distinctively, LazyEviction maintains the original output length of LLMs while enhancing memory and computational efficiency through optimized KV cache usage, demonstrating compatibility with these compression approaches.
System-level Optimizations. By optimizing GPU memory usage and batch processing, vLLM [33] and FlashDecoding [35] imporve both time and memory efficiency in decoding. Additionally, QUEST [21] and OmniKV [22] store KV caches in CPU and dynamically schedule them during the decoding phase. Our method is based on an eviction strategy that prioritizes the importance of tokens in the future, and it is parallel and compatible with the above methods.
@@ 3 Observations and Motivations
Figure 2: (a) shows the performance degradation for SOTA Methods on reasoning tasks. With the same KV cache compression ratio (e.g., 50%), compared with traditional language modeling tasks dataset PG-19 [23], the performance of both H2O and TOVA has decreased by ~20% on GSM8K dataset. (b) is the visualization of the importance variation by selecting Top-50% important tokens. Tokens at the same position show different importance at different decoding steps.
Figure 3: Visualization of TIR. We observe attention maps across different heads of DeepSeek-R1-Distill-Llama-8B. We find most tokens(>95%) show TIR pattern. (a) and (b) show an attention map with recurring tokens and their corresponding positions. (c) statistically analyzed the MRI distribution in different models among different tasks.
The goal of KV cache eviction is to identify and evict KV pairs that are non-critical for future decoding steps. To preserve the correctness of subsequent decoding results, it is critical to retain tokens with potential future importance. In this section, we first analyze the previous works and observe attention patterns in reasoning tasks.
Finding 1
Exiting works fail to reserve the potentially important tokens during decoding steps, thus exhibiting degraded performance in long reasoning tasks.
We compare the performance of H2O [16] and TOVA [13] on reasoning tasks versus general long-context generation tasks in Fig. 2(a), H2O and TOVA report competitive performance on standard language modeling dataset, but exhibit significant performance degradation on GSM8K under the same KV cache compression ratio. To diagnose this failure, we track the positions of top-50% important tokens across decoding steps in Fig. 2(b) and observe that tokens critical in later steps are often missing in earlier steps. Due to their low importance scores in intermediate steps, these tokens are prone to premature eviction in prior works.
Finding 2
Most tokens (>95%) exhibit recurrent attention patterns in multi-step reasoning tasks. Since they contain critical information for the reasoning, prematurely evicting may cause knowledge incontinuity.
@@
To further validate the generality of this phenomenon, we measured the Maximum Recurrence Interval (MRI) — the longest period between consecutive re-activations of a token — across multiple reasoning models on the GSM8K and MATH500 datasets. Results show that over 95% of tokens exhibit importance recurrence (MRI > 1), confirming its ubiquity in reasoning tasks (Fig. 3(c)).
Finding 3
Most recurring tokens’ MRI is far smaller than the output length and can be detected via an observation window.
Figure 4: Overview of our proposed LazyEviction Framework, where the dark squares represent that the token has a relatively higher attention score. (a) LazyEviction performs eviction decisions at intervals of $W$ steps. The workflow contains two key operations: (b) Dynamic MRI Tracking according to updated important timestamps, and (c) MRI-Centric Scoring during decision phases, where tokens predicted to be critical for future steps are retained. (d) The MRI-Centric Score fundamentally predicts future token importance by analyzing historical patterns of importance variation (i.e., MRI and the time elapsed since their latest timestamp).
As shown in Fig. 3(c), we statistically analyzed the MRI distribution in different models among different tasks. As the output length of the model increases, the MRI gradually becomes larger, but most tokens’ MRI remian relatively small. For example, the output length of Qwen model on the MATH500 dataset can reach 8k, but we have statistically found that 80% tokens’ MRI are less than 175. Therefore, as long as the window size exceeds the MRI of most tokens, the majority of recurring tokens will be identified. This finding motivates our design of the observation window-based lagged eviction mechanism.
@@ 4 LazyEviction
In this section, we introduce our LazyEviction, a novel framework of Observation Window-based Lagged Eviction. We first model LazyEviction as a multi-step dynamic optimization problem in Appendix D. Unlike existing works that perform KV eviction decisions at every decoding step, our LazyEviction adopts a window size $W$ as the interval for KV eviction decisions. As illustrated in Fig. 4(a), during decision-making, LazyEviction always retains the most recent $W$ KVs to preserve local coherence. When the number of KVs exceeds the predefined budget $B(B\gg W)$, LazyEviction executes KV eviction. The overall procedure of our LazyEviction is summarized in Appendix. B.
To select BWB-W KV pairs from past caches to retain, LazyEviction tracks importance variations (i.e., MRI) of each token at every decoding step (Recurrence Interval Tracking). For eviction decisions, importance scores, which are defined to indicate tokens’ potential future importance, are calculated based on MRI. Consequently, LazyEviction selectively retains only the tokens in the KV cache most likely to be critical in the future (MRI-Centric Eviction).
Recurrence Interval Tracking.
To identify importance variations for recurring tokens, we employ timestamps to track each token’s time of being important. Consistent with RaaS [19], whenever a token receives an attention score exceeding a threshold $\alpha$, its latest timestamp is updated to the value of current decoding step $t$. Thus, during the whole decoding phase, LazyEviction dynamically maintains a timestamp vector $\mathbf{TS}_{t}$ for all retained tokens, which records only the most recent activation time of tokens (illustrated in the leftmost figure in Fig. 4(b)).
However, only relying on $\mathbf{TS}_{t}$ fails to capture the temporal patterns of Importance Recurrence. To address this, we introduce MRI to record the longest interval between two consecutive activations in history. For newly generated tokens, MRI is initialized to 0. The MRI value is updated as:
$\mathbf{MRI}_{t}=\max\{\mathbf{MRI}_{t-1},\mathbf{TS}_{t}-\mathbf{TS}_{t-1}\}.$ (1)
MRI-Centric Eviction.
When the size of cached KV pairs $\left|{\mathcal{S}_{t}}\right|$ exceeds $B$, KV eviction are triggered at periodic intervals $t=kW$ ($k\in\mathbf{N}^{+}$). Leveraging the tracked historical importance dynamics via $\mathbf{MRI}_{t-1}$ and $\mathbf{TS}_{t}$. We aim to predict each token’s future importance using an MRI-Centric Importance Score. As illustrated in Fig 4(d), the importance score should be designed under the following two intuitive heuristics:
@@
    H1-score: Given a token $i$, if the time elapsed since its last activation ($t-T{S_{t}}[i]$) is greater than its $MR{I_{t}}[i]$, the possibility of token $i$ to be important in the future will decrease. The larger $\frac{{t-T{S_{t}}[i]}}{{MR{I_{t}}[i]}}$}, the less likely token $i$ is important. Thus, we define $\mathbf{H}^{(1)}_{t}[i]=2\sigma(-{\frac{{t-T{S_{t}}[i]}}{{MR{I_{t}}[i]}}})$, where $\sigma(x)=1/(1+\exp(-x))$ is sigmoid function.
@@
    H2-score: Tokens with smaller MRI values have a higher possibility to be important in the future, as frequently recurring tokens are deemed more critical. Thus, we define $\mathbf{H}^{(2)}_{t}[i]=2\sigma(-{{\frac{1}{{MR{I_{t}}[i]}-1}}})$. Notably, if $MR{I_{t}}[i]=0$, the token has never been activated since its generation (MRI is initialized as 0), resulting in $\mathbf{H}^{(2)}_{t}[i]=0,\rm{if}MR{I_{t}}[i]=0$.
H1-Score reflects the likelihood of the token regaining importance within the next observation window, while H2-Score prioritizes tokens with smaller MRI values. We discuss the importance of these two score functions in Sec 5.3, and explain the rationale behind the formulation of each function in Appendix D. Thus, the MRI-Centric Importance Score can be formalized as:
@@
For each eviction decision, we first compute $\mathbf{I}_{t}$ with Eq. 2, then retain the most recent $W$ and $\rm{Top}(B-W)$ KV pairs with the highest scores.
Notably, LazyEviction strategy introduces additional computational overhead, which may affect inference efficiency. We analyze this in Appendix E through both theoretical analysis and empirical evaluation. We point out that LazyEviction mitigates additional costs by making eviction decisions at fixed intervals rather than continuously. This $W$-steps eviction mechanism results in significantly lower overhead compared to existing KV compression methods. Moreover, since LazyEviction limits the KV budget during decoding steps, as the number of generated tokens increases, the computational efficiency does not grow linearly like it does with Full KV. When the generated tokens reach 16k, the overall inference efficiency of LazyEviction surpasses Full KV.
Disscussion about the choice of $W$ and $\alpha$.
In LazyEviction, the observation window with size $W$ serves to detect recurring tokens. However, since the most recent $W$-th tokens must be preserved, the setting of $W$ inherently introduces a trade-off between local and global information retention. An excessively large $W$ risks losing critical past tokens, while an overly small $W$ may fail to detect numerous recurring tokens. To resolve this dilemma, we employ the observation from Fig. 3(c), setting $W$ as the MRI threshold corresponding to the 80% tokens. This threshold can be statistically determined through offline analysis by randomly selecting 1% samples from the test datasets. Our experiments in Appendix F.1 analyze how $W$ affects reasoning performance.
Since the selection of $\alpha$ influences the temporal distribution of token timestamps, both extremely large and small $\alpha$ hinder effective identification of tokens’ importance variation. As the methodology for determining $\alpha$ has been thoroughly discussed in RaaS [19], we directly adopt their approach in LazyEviction. We analyze the effect of setting different $\alpha$ in Appendix F.2.